Twenty-nine (29) earthquakes were located during this period. Most of the events were located beneath the Soufriere Hills volcano, between 0 to 2 km depth (0 to 1.5 miles). One earthquake occurred at Garibaldi Hill. The level of seismicity remains quite high with tremor amplitudes increasing since 0400 today. The level of tremor at the volcano is now higher than during any other period of the current crisis.
EDM measurements were carried out at Whites and at the new site at Long Ground today. The new measurements will allow a more accurate determination of the deformation which has been occurring in the crater. The EDM instrument was also used to take angular measurements to the new spine from several locations around the volcano. The angles would be remeasured daily and would be used to help determine the rate at which the spine is growing.
Comparison of the GPS measurements made three days ago with those done previously indicate that there has been 1cm lengthening along all the lines to Harris' Lookout. Technical difficulties prevented any measurements from being made yesterday.
Excellent conditions today once again provided very good views of the crater area from the helicopter. The new spine close to Castle Peak, which was first observed on Sunday, continue to increase in height at a relatively rapid rate. Several cracks along the flanks of this feature have increased in size since last observed yesterday evening. The spine continue to appear increasingly unstable. Steam and light ash emission from the crater area increased during early afternoon and continued throughout the evening. Several distinct puffs of reddish ash was emitted and observed by scientist conducting EDM measurements during the late afternoon.